The bases of a trapezoid are 2.4 dm. and 5.6 dm. and the altitude is 5 dm. Find the

area of the trapezoid.



1
Expert's answer
2021-01-28T12:40:26-0500

"A=\frac{a+b}{2}h=\frac{2.4+5.6}{2}*5=20\;dm^2."